At maximum speed, an airplane travels 1720 miles against the wind in 5 hours. Flying with the wind, the airplane can travel that same distance in 4 hours. X= maximum speed of plane Y= speed of wind. What is the speed of the plane with no wind?
Answer by checkley71(8403)

1720/x=5
5x=1720
x=1720/5
x=344 mph against the wind.
1720/y=4
4y=1720
y=1720/4
y=430 mph. with the wind.
(430-344)/2
86/2=43 mph. is the speed of the wind.
344+43=387
430-43=387 mph. speed of the plane in still air.